OpenStudy (anonymous):

2^x+1, how would I draw the graph?

OpenStudy (solomonzelman):

x+1 is the exponent, or x is the exonent and 1 is added to the output?

OpenStudy (solomonzelman):

\(\large \color{black}{ 2^x+1 }\) or \(\large \color{black}{ 2^{x+1} }\) ?

OpenStudy (anonymous):

the first one

OpenStudy (solomonzelman):

So, you are just shifting it vertically. For example, \(\Large\color{ purple }{\Large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= a^x ~~~~~\rm{\Rightarrow}~~~~ g(x)= a^{\Large (x \color{red}{ -~\rm{c} })} &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= a^x ~~~~~\rm{\Rightarrow}~~~~ g(x)= a^{\Large (x \color{red}{ +~\rm{c} })} &~\rm{to~~the~~left~} \\ \text{ } \\ f(x)= a^x ~~~~~\rm{\Rightarrow}~~~~ g(x)= a^x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= a^x ~~~~~\rm{\Rightarrow}~~~~ g(x)= a^x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

sorry it took too long to type this.

OpenStudy (anonymous):

Do I do anything with the 2? @SolomonZelman

OpenStudy (solomonzelman):

Well, the 2^x is the essence of your graph. And this is what it looks like. https://www.desmos.com/calculator/35vkgex12p you can notice, that x, the domail, is all real numbers. Because there is no restriction on what to put in. Dom: \(\large\color{black}{ (-\infty~,~+\infty) }\) The y will approach zero, as x becomes an exponent of negative values with bigger magnitudes.

OpenStudy (solomonzelman):

you just need to observe the pattern and do an accurate, approximate, but not absolutely exact sketch. \(\large\color{black}{ y=2^x }\) \(\large\color{black}{ x=0,~~y=2^0~~\Rightarrow ~~~y=1~~~~~~\Rightarrow (0,1)}\)

OpenStudy (solomonzelman):

Do just a couple of points like this.

OpenStudy (solomonzelman):

Oh my man, this site is so glitchy!!

OpenStudy (solomonzelman):

\(\large\color{black}{ x=1,~~y=2^1~~\Rightarrow ~~~y=2~~~~~~\Rightarrow (1,2)}\) \(\large\color{black}{ x=2,~~y=2^2~~\Rightarrow ~~~y=4~~~~~~\Rightarrow (2,4)}\) \(\large\color{black}{ x=3,~~y=2^3~~\Rightarrow ~~~y=8~~~~~~\Rightarrow (3,8)}\)

OpenStudy (solomonzelman):

See the pattern for all positive values?

OpenStudy (solomonzelman):

I don't how you want to graph it. You can use a graphing calculator, or by hand, what exactly do you want?

OpenStudy (anonymous):

How would I do it if it was \[3^{x-1}\]

OpenStudy (anonymous):